OpenStudy (anonymous):

An obj. that is projected straight up with an initial velocity v feet per second and initial height s feet, t is time in seconds. trent throws a baseball straight upward with a velocity of 44 feet per second from a height of 12 feet.In how many seconds will the ball hit the ground?

OpenStudy (anonymous):

initial velocity u = 44ft/sec s= ut + 1/2 a t^2 -12 = 44t + (0.5)( -32.2) t^2 (assume a=32.2ft/sec) 16.1t^2 -44t -12 = 0 so t = { -(-44) + sqrt[(-44)^2 -4(16.1)(-12)]} / (2*16.) t=2.98 sec